# Dependency-Aware Semi-Structured Sparsity: Declining Roles of Outliers in Pruning GLU-based LLMs

### v2

@@ Discussion.
Table 2: WikiText perplexity of pruned LLMs. The (*) in Size indicates a model that uses larger MLP than Vallina Transformer (e.g., grouped query attention).
5 Experiments
5.1 Settings
Models.
DaSS’s performance is evaluated over open LLMs using GLU variants. SwiGLU is the most widely used GLU-based MLP in the recent LLMs, including the LLaMA2 model family Touvron et al. (2023), which has models with parameters ranging between 7 billion and 70 billion, and also the Mistral-7B model Jiang et al. (2023). Among them, LLaMA2-70B and Mistral-7B use grouped-query attention Ainslie et al. (2023), the MLP module parameter accounts for around 80% of the total model parameters. There are only a few open LLMs that use other variants. For GeGLU, we use Gemma-7B. It is worth noting that the MLP intermediate dimension of Gemma-7B is 8×8$\times$8 × model dimension, making the MLP module much larger. For ReGLU, we use ReluLLaMA Team (2023), which is fine-tuned using ReGLU variant Shazeer (2020); Mirzadeh et al. (2023) based on LLaMA2 with small accuracy loss. The model configuration details are in Appendix A.1. We access the public checkpoints of the involved models provided by HuggingFace Transformers Wolf et al. (2019).
Baseline Approaches.
We compare the performance with two LLM-specific one-shot pruning approaches, SparseGPT Frantar and Alistarh (2023) and Wanda Sun et al. (2024). We don’t consider structured pruning methods like LLM-Pruner Ma et al. (2023) and Sheared LLaMA Xia et al. (2023b), as they typically require re-training to recover accuracy, and are less practical for large models like LLaMA2-70B. Those baseline methods utilize uniform layerwise sparsity that can be easily converted into hardware-friendly N:M sparsity pattern. We used the same calibration data set as SparseGPT and Wanda in their model pruning processes, consisting of 128 sequences of 2048 tokens each, randomly selected from the first shard of the C4 dataset Raffel et al. (2020).
Evaluation.
To comprehensively evaluate the efficacy of our proposed method, two unique metrics are utilized to evaluate the performance of the pruned models: (1) perplexity (PPL) of language modeling (2) zero-shot accuracy on 5 commonsense reasoning tasks. Perplexity has been regarded as a consistent and reliable metric for measuring compressed models Dettmers and Zettlemoyer (2023); Frantar and Alistarh (2023), downstream tasks tend to be noisy, but more interpretable. For perplexity evaluation, we use the validation dataset of WikiText2 Merity et al. (2017). For zero-shot commonsense reasoning tasks, we choose five widely used tasks for accuracy evaluation: ARC (Easy and Challenge) Clark et al. (2018), HellaSwag Zellers et al. (2019), PiQA Bisk et al. (2020), and WinoGrande Sakaguchi et al. (2021), implemented in the Lm-Evaluation-Harness Gao et al. (2021). We evaluate the perplexity of all the aforementioned models. To fully demonstrate the task-wise performance in different sparsity patterns, we report the downstream task performance of the largest LLaMA2-70B model.
Sparsity.
In the less interpretable perplexity evaluation, we only prune the MLP layers. In the task-wise evaluation of LLaMA2-70B model, we prune both attention and MLP modules in accordance with previous works to better understand the performance gap between the pruned and original models. For attention module pruning, we use Wanda method, which is more efficient than SparseGPT. In LlaMA2-70B, as the MLP module accounts for more than 80% total parameters, we observe the choice of attention pruning has almost no impact on the final performance. We apply a uniform sparsity ratio across all the pruned layers and evaluate three sparsity types: unstructured sparsity, and semi-structured sparsities of 4:8 and 2:4.
@@ 5.2 Language Modeling
This evaluation involved a detailed examination of all the aforementioned LLMs. In perplexity evaluation, as shown in Table 2, our method consistently achieves better performance than SparseGPT and Wanda in more constrained and practical N:M sparsity pattern. As indicated by Sun et al. (2024), where weight updates can improve the performance in N:M sparsity pattern, our method shows superior performance even without computationally expensive weight updates. For Mistral-7B, Gamma-7B and LLaMA2-70B models with larger MLP layers, our method also outperforms SparseGPT in unstructured sparsity. Impressively, for the Gemma-7B with much larger MLP layers, DaSS outperforms Wanda in N:M sparsity significantly. DaSS shows consistent effectiveness across SwiGLU, GeGLU, and ReGLU, proving the generalizability of the DaSS method on GLU variants.
@@ 5.3 Downstream Tasks
While perplexity has been a popular metric in earlier LLM compression works Frantar and Alistarh (2023); Dettmers and Zettlemoyer (2023), its result is less interpretable. Apart from assessing perplexity, we comprehensively evaluate the performance of pruned LLaMA2-70B models in the widely used commonsense reasoning tasks.
In Table 3, we present the performance of different sparse LLaMA2-70B models on downstream tasks with prompting. The results show that our method outperforms SparseGPT and Wanda in most tasks at semi-structured N:M sparsity pattern. The only exception is that SparseGPT outperforms both Wanda and DaSS in Winogrande task. 50% unstructured SparseGPT pruned model even outperforms the dense model. Such results align with the prevailing LLM compression works Dettmers and Zettlemoyer (2023); Frantar and Alistarh (2023), where single task results can be noisy, the mean accuracy provides stable and reliable results. For unstructured sparsity, our method outperforms Wanda sharing the same complexity level.
@@ 5.4 Performance Analysis
Table 4 illustrates a comparison of the mean zero-shot task accuracy at varying levels of sparsity for the MLP module of the LLaMA2-70B model. It’s evident that DaSS pruning maintains competitive performance which closely matches that of SparseGPT across the entire range of sparsity ratios tested. Notably, DaSS achieves such performance without the need for weight updates, suggesting its effectiveness and efficiency in locating sparse neural networks. On the other hand, outlier-dependent Wanda pruning method shows a significant decline in accuracy as the sparsity ratio increases. This clear empirical evidence challenges the prevailing notion Dettmers et al. (2022); Lin et al. (2024) that weights corresponding to outliers are more important.
5.5 Running Time Analysis
Pruning speed
@@ 5.5 Running Time Analysis
Table 5: Pruning speed (in seconds) comparison.
@@
Table 6: Decode inference performance for varying sparsity levels with LLaMA2-7B utilizing DeepSparse NeuralMagic (2021) inference engine.
@@
We prune both attention and MLP module to test speed up. The speedup achieved by sparse model, as demonstrated in Table 6, highlights a significant reduction in end-to-end decode latency when utilizing LLaMA2-7B on the DeepSparse inference engine NeuralMagic (2021). The speed is tested on an Intel Xeon Platinum 8160 CPU with 24 cores. We highlight that the inference speedup is not exclusive to our pruning method but is a result of the inherent power of sparsity in accelerating computation.
@@ 6 Conclusion
Our observations reveal a diminishing importance of outliers in GLU-based MLP modules. Based on this insight, we propose the Dependency-aware Semi-structured Sparsity (DaSS) method. DaSS effectively balances the adaptability of unstructured pruning with the structural consistency of dependency-based structured pruning, utilizing MLP intermediate activation norms as the group importance indicator to develop a novel pruning metric. Empirical evaluations on the Mistral, Gemma and LLaMA2 model families show that DaSS surpasses state-of-the-art pruning methods such as SparseGPT and Wanda in achieving hardware-friendly N:M sparsity patterns. We hope our work inspires further research and development in the compression of GLU-based LLMs.
Limitations
Our research has the following limitations:
Merits beyond inference efficiency
Beyond improving inference efficiency, pruning techniques have demonstrated additional benefits for LLMs. Specifically, pruned models tend to exhibit fewer hallucinations compared to their unpruned counterparts Chrysostomou et al. (2024). Additionally, pruning has been shown to enhance resilience against jailbreak prompting Hasan et al. (2024) and to help in reducing toxicity and bias Xu and Hu (2022). However, the potential advantages of DaSS pruning beyond inference efficiency remain unclear and warrant further investigation.
Speed test in GPU
For LLM inference, the primary contributor to inference latency is the token generation process. While PyTorch’s current 2:4 sparsity implementation is optimized for accelerating matrix multiplication, it does not significantly enhance the performance of token generation. Recent work nm-vllm Shaw and Goin (2024) has shown approximately 1.5×1.5$\times$1.5 × token generation speedup for LlaMA2-7B at 50% unstructured sparsity. However, since this library is newly released and still under active development, it currently has many bugs, and we have not yet successfully implemented it. Kurtic et al. (2023) reports the speedup of several N:M sparsity patterns, its implementation hasn’t been released yet. The hardware and software support for sparsity is rapidly advancing, which will make sparse models more user-friendly in the future.
Application to other model architectures
Our method leverages a newly discovered property of GLU-based MLPs, making it unsuitable for older LLMs, such as OPT Zhang et al. (2022), which do not incorporate GLU variants. Additionally, our approach does not extend to the attention module, as the outlier activations still play an important role in the attention module.
@@ A.1 Model Configurations
A.2 Robustness to calibration samples.
Ashkboos et al. (2023) observes that the intermediate activations of SwiGLU-based MLP layers exhibit high variance, primarily caused by the Hadamard product of the preceding two outputs. This leads to diminished accuracy in 4-bit quantization. In Figure 2, we present how varying the number of sequences sampled for calibration affects the performance of pruning methods. Even though only using intermediate activations, our method demonstrates minimal sensitivity to changes in the number of calibration samples.
Figure 2: Robustness to calibration samples.
A.3 GeLU-based MLP Input projection
In the older MLP module, unlike GLU variants, there are only two projection matrices. We investigate magnitude pruning in different granularity for the MLP input projection matrix. We use the Falcon-7B model, which uses GeLU-based MLP. "Layer Magnitude" means we group weights in the entire weight matrix using magnitude pruning. In Table 7, we surprisingly find the input-balanced magnitude pruning achieves much worse results. This is in contrast to the results in GLU variants. The results show the huge significance of input activations for GeLU-based input projection pruning. Since our DaSS method needs to group weights per input, our method is less applicable to older MLP modules. However, GeLU-based MLP is rarely used in the current state-of-the-art LLMs.
Sparsity 50% 60% 70%
Wanda 7.17 8.72 18.95
Output-balanced Magnitude 8.46 15.27 498.77
Input-balanced Magnitude 1708 10587 13706
Layer Magnitude 21.82 1895 12896
Table 7: The Wikitext perplexity results of GeLU input projection at different sparsity ratio
A.4 Combination with Quantization
@@
Pruning and quantization, traditionally seen as distinct model compression techniques, are not mutually exclusive and can be effectively integrated to enhance overall efficiency Han et al. (2015a); Frantar and Alistarh (2023); Agarwalla et al. (2024). Here we test the combination of DaSS with 4-bit AWQ Lin et al. (2024) for compressing LlaMA2-7B. As shown in Table 8, the performance of all the pruned models just drops slightly after 4-bit quantization. DaSS still outperforms wanda after 4-bit quantization.
A.5 MMLU Results
Concurrent work (Gromov et al., 2024) indicates the unusual behavior of LLMs in performing Massive Multitask Language Understanding (MMLU) Hendrycks et al. (2021) tasks. We use Chain-of-Thought Hub Fu et al. (2023) which is based on the official implementation of MMLU Hendrycks et al. (2021). The MMLU encompasses 57 tasks, spanning from STEM, Humanities, Social Sciences, among others and we report the mean accuracy of 57 tasks. We test the performance of pruned LlaMA2-70B model in MMLU task.
@@ A.5 MMLU Results
Table 9: MMLU scores for different sparsity methods and levels
From Table 9, we observe that the improvement of DaSS over Wanda becomes more pronounced in the challenging MMLU task, where it achieves an increase in accuracy of 1.17 and 1.12 for the 4:8 and 2:4 sparsity patterns, respectively. In accordance with the observations with Jaiswal et al. (2023), we see a considerable performance degradation in knowledge-intensive MMLU task, in which only unstructured 50% sparsity models outperform the dense LLaMA2-34B model. As GPU plays a more important role in larger model inference, it is essential to improve the performance of pruned models in hardware-friendly N:M sparsity. Uniform reduction of the overall sparsity level is not feasible for N:M sparsity, Frantar and Alistarh (2023) suggests a specific subset of layers can be chosen for full N:M sparsification. Here we skip pruning 1/4 consecutive layers (20 layers) and result in a final 37.5% sparsity ratio. We use the first 10 tasks of MMLU to study pruning sensitivity. We divide the model into 4 consecutive parts and study skipping pruning each part. As shown in Table 10, the earlier layers are more sensitive than the later ones in knowledge-intensitive tasks, which is contradictory to the findings in Frantar and Alistarh (2023) using perplexity, and align with the findings of Gromov et al. (2024).
@@
Model Param Layers Hidden Intermediate Query Heads KV Head
LlaMa2-7B 7B 32 4096 11008 32 32
LlaMa2-13B 13B 40 5120 13824 40 40
LlaMa2-70B 70B 80 8192 28672 64 8
Mistral-7B 7B 32 4096 14336 32 8
Gemma-7B 7B 28 3072 24576 16 16
Table 11: Model configurations of Llama2, Gemma and Mistral models.
@@
Concurrent work Gromov et al. (2024) propose an effective method to remove half of the layers in LlaMA2-70B with small accuracy loss for MMLU tasks, suggesting the unusual behavior of LLMs in performing MMLU tasks. Their work is focused more on task-specific pruning, while our work emphasizes preserving the general ability of LLMs to perform various tasks. Our weight pruning method is fundamentally orthogonal to layer pruning in principle.

### v3

@@ Discussion.
4 Experiments
4.1 Settings
Models.
DaSS’s performance is evaluated over open LLMs using GLU variants. SwiGLU is the most widely used GLU-based MLP in the recent LLMs, including the LLaMA2 model family (Touvron et al., 2023), which has models with parameters ranging between 7 billion and 70 billion, and also the Mistral-7B model (Jiang et al., 2023). Among them, LLaMA2-70B and Mistral-7B use grouped-query attention (Ainslie et al., 2023), the MLP module parameter accounts for around 80% of the total model parameters. There are only a few open LLMs that use other variants. For GeGLU, we use Gemma-7B. It is worth noting that the MLP intermediate dimension of Gemma-7B is 8×8$\times$8 × model dimension, making the MLP module much larger. For ReGLU, we use ReluLLaMA (Team, 2023), which is fine-tuned using ReGLU variant (Shazeer, 2020; Mirzadeh et al., 2023) based on LLaMA2 with small accuracy loss. The model configuration details are in Appendix A.1. We access the public checkpoints of the involved models provided by HuggingFace Transformers (Wolf et al., 2019).
Baseline Approaches.
We compare the performance with two LLM-specific one-shot pruning approaches, SparseGPT (Frantar & Alistarh, 2023) and Wanda (Sun et al., 2024). We don’t consider structured pruning methods like LLM-Pruner (Ma et al., 2023) and Sheared LLaMA (Xia et al., 2023), as they typically require re-training to recover accuracy, and are less practical for large models like LLaMA2-70B. Those baseline methods utilize uniform layerwise sparsity that can be easily converted into hardware-friendly N:M sparsity pattern. We used the same calibration data set as SparseGPT and Wanda in their model pruning processes, consisting of 128 sequences of 2048 tokens each, randomly selected from the first shard of the C4 dataset (Raffel et al., 2020).
Evaluation.
To comprehensively evaluate the efficacy of our proposed method, two different metrics are utilized to evaluate the performance of the pruned models: (1) perplexity (PPL) of language modeling (2) zero-shot accuracy on 5 commonsense reasoning tasks. Perplexity has been regarded as a consistent and reliable metric for measuring compressed models (Dettmers & Zettlemoyer, 2023; Frantar & Alistarh, 2023), while downstream tasks sometimes have tendency in noisy behavior, but more interpretable. For perplexity evaluation, we use the validation dataset of WikiText2 (Merity et al., 2017). For zero-shot commonsense reasoning tasks, we choose five widely used tasks for accuracy evaluation: ARC (Easy and Challenge) (Clark et al., 2018), HellaSwag (Zellers et al., 2019), PiQA (Bisk et al., 2020), and WinoGrande (Sakaguchi et al., 2021), implemented in the Lm-Evaluation-Harness (Gao et al., 2021). We evaluate the perplexity of all the aforementioned models. To fully demonstrate the task-wise performance in different sparsity patterns, we report the downstream task performance of the largest LLaMA2-70B model. Notably, LLaMA2-70B utilizes grouped-query attention (Ainslie et al., 2023), and the MLP module accounts for more than 80% of the total parameters, representing the most widely adopted architectural design in modern LLMs.
Sparsity.
In the less interpretable perplexity evaluation, we only prune the MLP layers. In the task-wise evaluation of the LLaMA2-70B model, both attention and MLP modules are pruned, consistent with prior works to accurately assess the performance gap between pruned and original models. Since DaSS is not applicable to attention module pruning, we employ the Wanda method to prune the attention module, ensuring overall efficiency remains consistent with Wanda. We apply a uniform sparsity ratio across all the pruned layers and evaluate three sparsity types: unstructured sparsity, and semi-structured sparsities of 4:8 and 2:4.
Table 1: WikiText perplexity of pruned LLMs. Here we only prune the MLP module. The (*) in Size indicates a model that uses larger MLP than Vallina Transformer (e.g., grouped query attention).
@@ 4.2 Language Modeling
We examined all the aforementioned LLMs in terms of perplexity as shown in Table 1. Our method consistently achieves better performance than SparseGPT and Wanda in more constrained and practical N:M sparsity pattern. As indicated by Sun et al. (2024), where weight updates can improve the performance in N:M sparsity pattern, our method shows superior performance even without computationally expensive weight updates. For Mistral-7B, Gamma-7B and LLaMA2-70B models with larger MLP layers, our method also outperforms SparseGPT in unstructured sparsity. Impressively, for the Gemma-7B with much larger MLP layers, DaSS outperforms Wanda in N:M sparsity significantly. DaSS shows consistent effectiveness across SwiGLU, GeGLU, and ReGLU, proving the generalizability of the DaSS method on GLU variants.
@@ 4.3 Downstream Tasks
Apart from assessing perplexity, we comprehensively evaluate the performance of pruned LLaMA2-70B models in the widely used commonsense reasoning tasks.
In Table 2, we present the performance of different sparse LLaMA2-70B models on downstream tasks with prompting. The results show that our method outperforms SparseGPT and Wanda in most tasks at semi-structured N:M sparsity pattern. The only exception is that SparseGPT outperforms both Wanda and DaSS in Winogrande task. 50% unstructured SparseGPT pruned model even outperforms the dense model. Such results align with the prevailing LLM compression works (Dettmers & Zettlemoyer, 2023; Frantar & Alistarh, 2023), where single task results can be noisy and the mean accuracy provides stable and reliable results. For unstructured sparsity, our method outperforms Wanda sharing the same complexity level.
@@ 4.4 Performance Analysis
Table 3 illustrates a comparison of the mean zero-shot task accuracy at varying levels of sparsity for the MLP module of the LLaMA2-70B model. It’s evident that DaSS pruning maintains competitive performance which closely matches that of SparseGPT across the entire range of sparsity ratios tested. Notably, DaSS achieves such performance without the need for weight updates, suggesting its effectiveness and efficiency in locating sparse neural networks. On the other hand, output-balanced Wanda pruning shows a significant decline in accuracy as the sparsity ratio increases. This suggests that Wanda pruning may suffer from structural mismatch issues within the MLP layer, which become more pronounced at higher sparsity levels. As a result, the neural network’s performance deteriorates, potentially leading to a dysfunctional model at extreme sparsity ratios.
Robustness to calibration samples.
Ashkboos et al. (2023) observes that the intermediate activations of SwiGLU-based MLP layers exhibit high variance, primarily caused by the Hadamard product of the preceding two outputs. This leads to diminished accuracy in 4-bit activation quantization. In Figure 2, we present how varying the number of sequences sampled for calibration affects the performance of pruning methods. Even though only using intermediate activations, our method demonstrates minimal sensitivity to changes in the number of calibration samples.
Figure 2: Robustness to calibration samples.
Ablation Study on Hyperparameter $\alpha$
We conducted an additional ablation study to investigate the impact of different values of the group importance strength hyperparameter $\alpha$ on the performance of DaSS. Specifically, we experimented with LLaMA2-7B using $\alpha=\{0.25,0.5,0.75,1.0\}$. The results are presented in Table 5.
The ablation results show that $\alpha=0.25$ achieves slightly better performance compared to $\alpha=0.5$ for LLaMA2-7B in terms of WikiText perplexity. However, we note that $\alpha=1.0$ still outperforms Wanda (PPL: 6.50). We used $\alpha=0.5$ for all experiments without intentionally tuning this hyperparameter for different models and tasks.
Table 4: WikiText perplexity of 50% MLP sparsity for LLaMA2-7B using different values of $\alpha$.
$\mathbf{\alpha}$ PPL (↓)
1.0 6.48
0.75 6.46
0.5 6.44
0.25 6.43
Table 5: Pruning speed (in seconds) comparison.
@@
4.5 Running Time Analysis
Pruning speed
@@ 4.5 Running Time Analysis
Table 6: Decode throughput performance for varying sparsity levels with LLaMA2-7B utilizing DeepSparse NeuralMagic (2021) inference engine.
@@
We prune both attention and MLP module to test the inference speed. The speedup achieved by sparse model, as demonstrated in Table 6, highlights a significant reduction in end-to-end decode latency when utilizing LLaMA2-7B on the DeepSparse inference engine (NeuralMagic, 2021). The speed is tested on an Intel Xeon Platinum 8160 CPU with 24 cores. We highlight that the inference speedup is not exclusive to our pruning method but is a result of the inherent power of sparsity in accelerating computation.
5 Related Work
Pruning LLM. Neural network pruning in LLM can be broadly categorized into two groups: structured pruning (Ma et al., 2023; Zhang et al., 2023) and unstructured pruning (Frantar & Alistarh, 2023; Sun et al., 2024). Ma et al. (2023) proposes a dependency detection algorithm to detect and prune non-critical grouped structures followed by LoRA fine-tuning. Although structured pruning can usually achieve better hardware efficiency, the accuracy drops a lot even at a low compression rate. Unstructured pruning can yield a higher compression rate and achieve acceleration on Nvidia’s GPUs by employing a hardware-friendly N:M sparsity pattern. SparseGPT (Frantar & Alistarh, 2023) leverages the Hessian inverse for pruning and reduces reconstruction error of dense and sparse weights by subsequent weight updates. Wanda (Sun et al., 2024) employs an efficient method that augments input activations into weight magnitudes, and matches the performance of SparseGPT at medium sparsity. Our work incorporates dependency information into unstructured pruning, achieving a novel pruning paradigm.
Inherent Sparsity of Transformer MLP. Interestingly, sparsity within the MLP activations of trained Transformer-based models occurs innately even without applying explicit regularizations or constraints (Zhang et al., 2022; Li et al., 2023; Dong et al., 2023). Such a phenomenon is prevalent in learned Transformers, including other zero-saturating functions. Liu et al. (2023); Mirzadeh et al. (2023); Zhang et al. (2022) achieve actual LLM inference speedup by only performing computation corresponding to the activating neuron for a given input. They do not actually reduce the model size since they mainly reduce I/O and computation latency in a selective weights loading manner, and thus, these methods are less applicable in large batch-size inference settings. Our work investigates the weight sparsity in MLP module by considering corresponding intermediate activations.
Outlier-dependent LLM Compression.
Outlier features, defined as features with magnitudes substantially larger than others, are a notable characteristic of LLMs (Dettmers et al., 2022). Despite making up only a small fraction of all feature dimensions, these outliers play a critical role in attention and predictive performance. Such observation has motivated the development of LLM-specific quantization methods (Dettmers et al., 2022; Xiao et al., 2023; Lin et al., 2024; Ashkboos et al., 2023) to handle outliers more effectively. Wanda (Sun et al., 2024) and OWL (Yin et al., 2024) broaden these insights, revealing that outlier features are significant in deciding weight importance when pruning LLMs. Our method diverges from conventional wisdom by demonstrating that, in the context of GLU-based MLP input projections, the inportance of input activation outliers is not as pronounced as previously assumed, prompting a reevaluation of their role in LLM pruning strategies.
@@ 6 Conclusion
We propose the Dependency-aware Semi-structured Sparsity (DaSS) method which effectively addresses the challenges of pruning GLU-based MLP modules LLMs. DaSS strikes a unique balance between the adaptability of unstructured pruning and the orderliness of structured pruning. By leveraging the MLP intermediate activation norms as the group importance indicator, we develop a novel pruning metric that assesses weight importance in a more structurally consistent manner. Empirical evaluations on the Mistral, Gemma and LLaMA2 model families show that DaSS surpasses state-of-the-art pruning methods such as SparseGPT and Wanda in achieving hardware-friendly N:M sparsity patterns. We hope our work inspires further research and development in the compression of GLU-based LLMs.
@@ A.1 Model Configurations
Table 7: Model configurations of Llama2, Gemma and Mistral models.
Model Param Layers Hidden Intermediate Query Heads KV Head
LlaMa2-7B 7B 32 4096 11008 32 32
LlaMa2-13B 13B 40 5120 13824 40 40
LlaMa2-70B 70B 80 8192 28672 64 8
Mistral-7B 7B 32 4096 14336 32 8
Gemma-7B 7B 28 3072 24576 16 16
A.2 Combination with Quantization
@@
Pruning and quantization, traditionally seen as distinct model compression techniques, are not mutually exclusive and can be effectively integrated to enhance overall efficiency (Han et al., 2015a; Frantar & Alistarh, 2023; Agarwalla et al., 2024). Here we test the combination of DaSS with 4-bit AWQ (Lin et al., 2024) for compressing LlaMA2-7B. As shown in Table 8, the performance of all the pruned models just drops slightly after 4-bit quantization. DaSS still outperforms wanda after 4-bit quantization.
A.3 MMLU Results
Recent work (Gromov et al., 2024) indicates the unusual behavior of LLMs in performing Massive Multitask Language Understanding (MMLU) (Hendrycks et al., 2021) tasks. We use Chain-of-Thought Hub (Fu et al., 2023) which is based on the official implementation of MMLU (Hendrycks et al., 2021). The MMLU encompasses 57 tasks, spanning from STEM, Humanities, Social Sciences, among others and we report the mean accuracy of 57 tasks. We test the performance of pruned LlaMA2-70B model in MMLU task.
Table 9: MMLU scores for different sparsity methods and levels
@@ A.3 MMLU Results
From Table 9, we observe that the improvement of DaSS over Wanda becomes more pronounced in the challenging MMLU task, where it achieves an increase in accuracy of 1.17 and 1.12 for the 4:8 and 2:4 sparsity patterns, respectively. In accordance with the observations with Jaiswal et al. (2023), we see a considerable performance degradation in knowledge-intensive MMLU task, in which only unstructured 50% sparsity models outperform the dense LLaMA2-34B model. As GPU plays a more important role in larger model inference, it is essential to improve the performance of pruned models in hardware-friendly N:M sparsity. Uniform reduction of the overall sparsity level is not feasible for N:M sparsity, Frantar & Alistarh (2023) suggests a specific subset of layers can be chosen for full N:M sparsification. Here we skip pruning 1/4 consecutive layers (20 layers) and result in a final 37.5% sparsity ratio. We use the first 10 tasks of MMLU to study pruning sensitivity. We divide the model into 4 consecutive parts and study skipping pruning each part. As shown in Table 10, the earlier layers are more sensitive than the later ones in knowledge-intensitive tasks, which is contradictory to the findings in Frantar & Alistarh (2023) using perplexity, and align with the findings of Gromov et al. (2024).
@@
Recent work (Gromov et al., 2024) propose an effective method to remove around half of the layers in LlaMA2-70B with small accuracy loss for MMLU tasks, suggesting the unusual behavior of LLMs in performing MMLU tasks. Their work is focused more on task-specific pruning, while our work emphasizes preserving the general ability of LLMs to perform various tasks. Our weight pruning method is fundamentally orthogonal to layer pruning in principle.